LIN AN a)2{sin 210 (1 - ) = HE b) & ( 33-25+1?
Answer #1

[{1+73,212 +27+ g = fue=],7 +13), 14 (14,3)+?], 12 { 0 13,1 = {1*575 3,764 2017315. {rn (1-1/8.37 = 5(2-1)N[(4-9 2 + 1] 445